How do you solve for x in the following equation?
log x + log(x+30)=3

Answer by scott8148(6628)

x(x + 30) = 1000

x^2 + 30x - 1000 = 0

(x + 50)(x - 20) = 0


remember, logarithms are not defined for values ≤0